Complete the proof:
Given: 4x + 8 = 2x - 12
Prove: x = -10

Answer by ikleyn(52879)

.

            It is about solving simple linear equation in one unknown --- one of simplest problems in Algebra.


Step 1.  Subtract 2x from both sides.  You will get then

             2x + 8 = -12.


Step 2.  Subtract 8 from both sides.  You will get

             2x = -12 - 8,  or

             2x = - 20.


Step 3.  In the last equation, divide both sides by 2.  You will get

             x = -20/2 = -10.


It is what the problem asks to prove.

The solution / (the proof) is completed.

--------------

To learn the subject  (on solving simple linear equations),  see the lesson
    - HOW TO solve a linear equation
in this site.

This subject is one of the basic Algebra skills.